# CO3: Contrasting Concepts Compose Better

### v1

@@ 4.2 Experimental Results
Comparison with Optimization-Free Approaches: To further demonstrate CO3’s model-agnostic nature and generalization to other base models, we evaluate it on PixArt-$\Sigma$ (Chen et al., 2024). We compare against both the base model and the sampling-based, optimization-free Composable Diffusion (Liu et al., 2022). Unlike Composable Diffusion, which composes diffusion chains within the DDIM prediction step, CO3 applies composition as a corrector. As shown in Fig. 5, Composable Diffusion suffers from missing or mixing concepts, while CO3 consistently preserves all concepts with correct bindings in the output.
Figure 5: Model Agnostic behavior: Qualitative comparison of generation from PixART-$\Sigma$(Chen et al., 2024) base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion.
@@ 5 Related Work
Optimization-based correction approaches: (1) Text-embedding optimization works like T2I-Zero variants (Tunanyan et al., 2023), and style or prompt inversion (Li et al., 2023a), improve semantic alignment by optimizing or restructuring the text representation during inference. These methods tweak or invert token embeddings so that the denoising trajectory better reflects entity–attribute bindings or multi-concept prompts.
@@
Composable generation works: view conditional diffusion models as energy/score functions, enabling algebraic composition. Composed-Diffusion (Liu et al., 2022) frames score-composition under CFG and demonstrates test-time generalization, but suffers from concept mixing and missing. Follow-ups propose training-free and model-agnostic methods like energy-parameterized diffusion and Metropolis/MCMC-corrected samplers that markedly improve multi-condition generation (Du et al., 2024), but their performance is often poor (Chefer et al., 2023; Feng et al., 2022). Kwon & Ye (2025) uses a composition similar to our resampler but not in the context of a corrector. Instead they use multiple DDIM forward-backward steps to sample initial noise.
@@ 6 Conclusion & Future Work
While our corrector’s closeness-aware weight modulation suppresses the concept mode overlap, the noise/score landscape is strongly biased by the training paradigms of diffusion models, i.e., the quantity and quality of the multi-concept bindings in the training set. This problem also manifests when presented with more unrealistic prompts, resulting in poor text/concept alignment. As shown in Fig. 7, while our work offers a significant boost to multi-concept compositional generation, clearly, there are failure cases that are not yet addressed. However, we believe that more advanced energy-based compositional samplers—that explicitly take into account the probabilities of the landscape during sampling—can further help in avoiding such cases. We hope to investigate these special cases in the future. We also hope to explore applications beyond composition using the CO3 corrector.
7 Reproducibility Statement
@@ A.2 Composition in Tweedie-denoised space for a valid cfg
Let’s define CFG as a function at time $t$ as, $f_{CFG}({\epsilon};\lambda)=$ ${\epsilon}^{\phi}+\lambda({\epsilon}-{\epsilon}^{\phi})$. Then for any $\lambda$ and $K>1$, $\sum_{k}{w_{k}f_{CFG}}=f_{CFG}(\sum_{k}{w_{k}{\epsilon}^{k}})$ only if $\sum_{k}w_{k}=1$.
@@
Let, $\tilde{{\epsilon}}^{\lambda}_{t}=\sum_{k}{w_{k}f_{CFG}}$. Then,
$\tilde{{\epsilon}}^{\lambda}_{t}$ $=\sum_{k}w_{k}({\epsilon}^{\phi}+\lambda({\epsilon}^{k}-{\epsilon}^{\phi}))$
$=\sum_{k}w_{k}{\epsilon}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-\sum_{k}w_{k}{\epsilon}^{\phi})$
$={\epsilon}_{\phi}\sum_{k}w_{k}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-{\epsilon}^{\phi}\sum_{k}w_{k})$
$={\epsilon}_{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-{\epsilon}^{\phi})\quad\text{if }\sum_{k}w_{k}=1$
$=f_{CFG}(\sum_{k}w_{k}{\epsilon}^{k})\quad\text{if }\sum_{k}w_{k}=1$
If $\sum_{k}w_{k}\neq 1$, then $\tilde{{\epsilon}}^{\lambda}_{t}\neq f_{CFG}(\sum_{k}w_{k}{\epsilon}^{k})$. ∎
@@
Lemma 3.
Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$. Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$. Then,
@@
    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
@@
$=x_{t}-\sigma_{t}\ f_{CFG}(\sum_{k}w_{k}{\epsilon}^{c_{k}})\quad\text{if }\sum_{k}w_{k}=1$
@@
A.3 More Implementation Details
A.3.1 CO3 Implementation Details
A.3.2 Baseline Methods
Attend-Excite (Chefer et al., 2023), Divide-Bind (Liu et al., 2023), InitNo (Guo et al., 2024), SynGen (Rassin et al., 2023) and Composable Diffusion (Liu et al., 2022) are methods based on SD1.5. We run experiments on these models using their publicly available code. ToMe (taihang Hu et al., 2024) is SDXL based we use their publicly available code-base implementation. We also adapted Composable Diffusion and SynGen to SDXL.
A.4 Time Complexity Analysis
We conduct an experiment to analyze the time-complexity of each SDXL based methods. The Table 3 indicates the average time (in secs) taken by the model per sample. The experiments are conducted on NVIDIA A100 gpu.
Method Inference Steps Time Cost (sec) Animals Animals-Objects Objects
SDXL 50 7.20 0.6950 0.8654 0.4926
Composable Diffusion 50 9.80 0.2846 0.5656 0.4529
SynGen(SDXL) 50 11.48 0.6816 0.8578 0.4652
ToMe 50 16.58 0.6257 0.8808 0.6440
CO3(ours) 50 19.9 0.7441 0.8878 0.5146
Table 3: Comparison of different methods across efficiency and compositionality metrics.
A.5 Comparison of compositions in score space
@@
Contrast this with the Composable-Diffusion’s noise/score composition:
@@
For arbitrary weights $\lambda_{k}$}, this cannot be expressed in the from equation 23. This proves that $\tilde{\epsilon}^{\lambda,C}_{t,compdiff}$ doesn’t lead to a valid Tweedie-mean.
A.6 More Visualizations
Figure 6: Qualitative results of SDXL base diffusion model, and SDXL + CO3 for animal-animal,animal-object, and object-object prompt categories
A.6.1 CO3 Failure Cases
A.7 Ablations on Hyperparameters

### v2

@@ 4.2 Experimental Results
ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$
Method A-A A-O O-O A-A A-O O-O
SDXL 0.7820 1.5474 0.6789 0.6951 0.8654 0.4926 0.8435
+ Resampling 1.0881 1.6755 0.8263 0.7351 0.8666 0.4528 0.9437
+ Corrector 1.0630 1.6429 0.8949 0.7177 0.8794 0.4796 0.9463
+ weight-mod. 1.2342 1.6744 1.0158 0.7441 0.8878 0.5146 1.0118
SDXL+Corrector 0.9543 1.5140 0.8894 0.7149 0.8801 0.4989 0.9086
+ weight-mod. 1.1732 1.5865 1.0067 0.7211 0.8701 0.5206 0.9797
Figure 5: Qualitative results of CO3 with different configurations on SDXL.
Qualitative Comparison: Figure 3 compares CO3 on Attend-Excite (simpler) prompts with several baselines. We illustrate common issues in compositional alignment—concept missing, attribute mixing, and object binding—across all categories. Divide-Bind (Liu et al., 2023) and Attend-Excite (Chefer et al., 2023) enhance the attention layer of the base model, but often yield cropped or incomplete concepts (rows 1, 3). Comp-Diff (Liu et al., 2022) suffers from missing or merged concepts (rows 2, 3). SynGen (Rassin et al., 2023) and ToMe (taihang Hu et al., 2024) perform well on Object prompts of the form “a [color1] [object1] and a [color2] [object2],” but still exhibit overlapping or mixed concepts (rows 2, 3; columns 5, 6). In contrast, CO3 achieves stronger object binding in the Animals-Animals and Animals–Objects categories while also improving results on Objects-Objects.
Figure 4 extends qualitative results to complex prompts from the Complex category of T2ICompbench and rare-concept prompts from Rarebench Park et al. (2024), comparing CO3 with the top-performing SDXL models (from Table 1). The left half shows prompts from T2ICompbench’s Complex category, which involves complex subject-attribute & inter-concept relations. Only CO3 is able to faithfully generate all concepts and relations, while the other methods often miss concepts or fail to capture relations. The right half shows prompts from RareBench, which involve rare or novel concepts. CO3 is able to generate extremely rare concepts such as ”banana shaped bottle” or ”Knighted turtle”, while the other methods often suffer from training data-bias – generating a regular banana or human knight instead.
Ablations Studies are preformed on the key components of CO3, starting from the base SDXL model (Table 3 & Figure 5). Adding resampling in the first three DDIM steps substantially improves average performance, particularly for the Animals and Animal–Objects categories. Since image layout is largely determined in the early timesteps, CO3-resampler iteratively explores better initial noise configurations that align with the prompt.
Figure 6: Model Agnostic behavior: Qualitative comparison of generation from PixART-$\Sigma$(Chen et al., 2024) base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion (Comp-diff).
Incorporating CO3-corrector over the next steps further improves prompts requiring fine-grained details such as color or texture, boosting performance on the Objects category in both ImageReward and BLIP-VQA.
Table 4: Generalization of CO3 on PixArt and comparison with model-agnostic baselines. cDiff denotes Comp-diff (Liu et al., 2022).
ImageReward $\uparrow$ BLIP-VQA $\uparrow$
Method A-A A-O O-O A-A A-O O-O
PixArt 1.066 1.697 1.023 0.691 0.861 0.559
PixArt + cDiff 0.917 0.047 0.339 0.208 0.591 0.472
PixArt + CO3 1.380 1.770 1.081 0.746 0.864 0.581
Additional gains come from the Closeness-Aware Concept Weight Modulation strategy, which consistently benefits all categories. We hypothesize that as the diffusion trajectory evolves, the latent state often drifts closer to one concept mode $c_{i}$; the modulation mechanism assigns $c_{i}$ a more negative weight, repelling the latent and mitigating concept dominance (Tunanyan et al., 2023) commonly observed in stable-diffusion.
Including the Corrector alone already improves performance in categories where attributes must attach to specific objects. This is most clearly visible in Rows 4–5 of the ablation table: prompts of the form “an [attribute1][object1] and [attribute2][object2]” show large gains, because early-timestep dominance (one object absorbing both attributes) is the failure mode that the Corrector resolves. Importantly, the average performance of SDXL + Corrector + weight-modulation (last row) is very close to SDXL + Resampling + Corrector + weight-modulation (Row 4), which indicates that the Corrector alone is powerful even without Resampling. However, the full improvement is achieved when the Corrector is applied after the Resampling steps.
Comparison with Optimization-Free Approaches: To further demonstrate CO3’s model-agnostic nature and generalization to other base models, we evaluate it on PixArt-$\Sigma$ (Chen et al., 2024) (Table 4). We compare against both the base model and the sampling-based, optimization-free Composable Diffusion (Comp-diff) (Liu et al., 2022). Unlike Comp-diff, which composes diffusion chains within the DDIM prediction step, CO3 applies composition as a corrector. As shown in Fig. 6, Comp-diff suffers from missing or mixing concepts, while CO3 consistently preserves all concepts with correct bindings in the output.
@@ 5 Related Work
Optimization-based correction approaches: (1) Text-embedding optimization works like T2I-Zero variants (Tunanyan et al., 2023; Zhuang et al., 2024), and style or prompt inversion (Li et al., 2023a), improve semantic alignment by optimizing or restructuring the text representation during inference. These methods tweak or invert token embeddings so that the denoising trajectory better reflects entity–attribute bindings or multi-concept prompts.
@@
Composable generation works: These works view conditional diffusion models as energy/score functions, enabling algebraic composition. Composed-Diffusion (Liu et al., 2022) frames score-composition under CFG and demonstrates test-time generalization, but suffers from concept mixing and missing. Follow-ups propose training-free and model-agnostic methods like energy-parameterized diffusion and Metropolis/MCMC-corrected samplers that markedly improve multi-condition generation (Du et al., 2024), but their performance is often poor (Chefer et al., 2023; Feng et al., 2022). On a different line, while Skreta et al. (2025) proposes a density estimation approach for composing diffusion chains for concept interpolation, Park et al. (2024) interpolates distributions of frequent and rare-concepts for improved compositional generation. Kwon and Ye (2025) uses a composition similar to our resampler but not in the context of a corrector. Instead they use multiple DDIM forward-backward steps to sample initial noise.
@@ 6 Conclusion & Future Work
While our corrector suppresses concept overlap, the model’s predicted score landscape remains biased by training data quality, occasionally leading to poor alignment in unrealistic prompts (Fig. 8). While our work offers a significant boost to multi-concept compositional generation, failure cases still persist and we believe advanced energy-based compositional samplers that explicitly model landscape probabilities can further mitigate these issues. We intend to investigate these edge cases and explore applications beyond compositional generation in future work.
Acknowledgments
Acknowledgments: This work was partially supported by NSF #2008338, #1909568, #2148583, and #MRI-2018966. This work used DELTA at NCSA through allocation CIS230230 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296
Reproducibility Statement
@@ A.2 Composition in Tweedie-denoised space for a valid cfg
Let’s define CFG as a function at time $t$ as, $f_{CFG}({\epsilon}_{t};\lambda)=$ ${\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}-{\epsilon}_{t}^{\phi})$. Then for any $\lambda$ and $K>1$, $\sum_{k}{w_{k}f_{CFG}({\epsilon}_{t}^{k};\lambda)}=f_{CFG}(\sum_{k}{w_{k}{\epsilon}_{t}^{k}};\lambda)$ only if $\sum_{k}w_{k}=1$.
@@
Let, $\tilde{{\epsilon}}^{\lambda}_{t}=\sum_{k}{w_{k}f_{CFG}({\epsilon}_{t}^{k};\lambda)}$. Then,
$\tilde{{\epsilon}}^{\lambda}_{t}$ $=\sum_{k}w_{k}({\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{k}-{\epsilon}_{t}^{\phi}))$
$=\sum_{k}w_{k}{\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{k}-\sum_{k}w_{k}{\epsilon}_{t}^{\phi})$
$={\epsilon}_{t}^{\phi}\sum_{k}w_{k}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{k}-{\epsilon}_{t}^{\phi}\sum_{k}w_{k})$
$={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{k}-{\epsilon}_{t}^{\phi})\quad\text{if }\sum_{k}w_{k}=1$
$=f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{k};\lambda)\quad\text{if }\sum_{k}w_{k}=1$
If $\sum_{k}w_{k}\neq 1$, then $\tilde{{\epsilon}}^{\lambda}_{t}\neq f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{k};\lambda)$. ∎
@@
Proposition 2.
(Restated) Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$. Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$. Then,
@@
    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
@@
$=x_{t}-\sigma_{t}\ f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}})\quad\text{if }\sum_{k}w_{k}=1$
@@
A.3 Comparison of compositions in score space
@@
Thus, when mixing scores from different conditional distributions, the above condition need to be satisfied for $\tilde{x}_{tweedie}$ to correspond to any coherent joint conditional model. Contrast this with the Composable-Diffusion’s (Liu et al. (2022)) noise/score composition:
@@
The above composition allows arbitrary weights, which in general violate the unit-sum constraint. In such cases, the result is simply a heuristic linear combination of scores, and the resultant Tweedie mean isn’t valid and doesn’t correspond to any well-defined distribution under the given forward process.
By contrast, CO3 composes Tweedie-means from different conditional distributions. Lemma 1 shows that these weights have to satisfy the same unit-sum condition as above for the resultant composed Tweedie-mean to be a valid one. CO3 maintains this condition and this preserves the update on the Tweedie manifold of the underlying sampler.
A.4 More Implementation Details
A.4.1 CO3 Implementation Details
A.4.2 Baseline Methods
Attend-Excite (Chefer et al., 2023), Divide-Bind (Liu et al., 2023), InitNo (Guo et al., 2024), SynGen (Rassin et al., 2023) and Composable Diffusion (Liu et al., 2022) are methods based on SD1.5. We run experiments on these models using their publicly available code. Tweediemix (Kwon and Ye, 2025), Magnet Zhuang et al. (2024), R2F (Park et al., 2024) and ToMe (taihang Hu et al., 2024) are SDXL based we use their publicly available code-base implementation. We also adapted Composable Diffusion and SynGen to SDXL.
Table 5: Human Evaluation results on Attend-Excite prompts. Higher the score, the better.
Method Animals Animals- Objects Objects
SDXL 0.5266 0.6050 0.7083
SynGen 0.3667 0.4366 0.7183
Magnet 0.3700 0.5683 0.6849
ToMe 0.4383 0.6083 0.6166
CO3(ours) 0.8817 0.7466 0.8250
A.4.3 Additional Human Evaluation
We conducted new human-evaluation studies (with 11 participants) across all benchmarks considered in this paper. For evaluation, we follow the protocol from R2F Park et al. (2024). For each prompt, images from CO3 and baseline methods are displayed side by side to enable direct comparison. Method names are anonymized and their order is randomly shuffled to reduce bias. Participants provide ratings in $\{0,1,2,3,4,5\}$, which are then normalized to $[0,1]$. Results are shown below (higher is better). Results on T2ICompbench and RareBench are included in Table 2 of the main text. Table 5 shows results on Attend-Excite prompts.
A.5 Generalization to higher-order solvers and low inference steps regimes
In addition to the main 50-step SDXL with DDIM experiments, we include results using both (i) low-step DDIM and (ii) low-step high-order DPM++ solvers (10–20 steps).
Please note that, since DDIM sampling has direct connections with denoising and generation via the Tweedie mean (Section 2 in the main paper), CO3 applies naturally to DDIM. However, our experiments below show that CO3 continues to provide consistent improvements under higher-order solvers and aggressive sampling budgets as shown in Table 6 & Table 7.
Table 6: Performance of CO3 with DPM++ solver and in the low inference steps regime.
Steps ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$
Method Animals Animals Objects Objects Animals Animals Objects Objects
SD-DDIM 50 0.7820 1.5574 0.6789 0.6950 0.8658 0.4925 0.8452
CO3-DDIM 50 1.2341 1.6743 1.0158 0.7441 0.8878 0.5146 1.0112
SD-DPM++ 20 0.7089 1.5652 0.6757 0.6974 0.8537 0.4984 0.8332
CO3-DPM++ 20 1.1424 1.5249 0.9537 0.7382 0.8842 0.4944 0.9563
SD-DPM++ 10 0.6834 1.5074 0.6681 0.6768 0.8537 0.4974 0.8145
CO3-DPM++ 10 1.0371 1.4520 0.9206 0.7331 0.8765 0.4443 0.9106
Table 7: Performance of CO3 with DDIM solver in the low inference steps regime.
ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$
Method Steps Animals Animals Objects Objects Animals Animals Objects Objects
SD w/ DDIM 50 0.7820 1.5574 0.6789 0.6950 0.8658 0.4926 0.8452
CO3 w/ DDIM 50 1.2341 1.6743 1.0158 0.7441 0.8878 0.5146 1.0112
SD w/ DDIM 20 0.6965 1.4567 0.5516 0.6832 0.8620 0.4947 0.7907
CO3 w/ DDIM 20 0.8532 1.6440 0.7481 0.6698 0.8611 0.4315 0.8671
SD w/ DDIM 10 0.4906 1.1503 0.1131 0.6428 0.8378 0.4752 0.6183
CO3 w/ DDIM 10 0.5872 1.5900 0.7117 0.6294 0.8589 0.4453 0.8037
A.6 PROMPTS WITH AND WITHOUT BACKGROUNDS
Figure 7: Concept obedient behavior of CO3: CO3 tends to strongly obey the concepts specified in the prompt; when the background is not specified it is conservative with hallucinating the background.
CO3 is strongly concept obedient. Many CO3 generated images from Attend-Excite prompts seem to lack background details. To investigate this behavior, we added background concepts to the same prompts shown in Figure 6 of the main text. Figure 7 shows the resulated images from this experiment. We observe that, CO3 tends to strongly obey the concepts specified in the prompt; when the background is not specified it is conservative with hallucinating the background. However, if the background is explicitly specified, CO3 brings the same degree of richness and quality to both the foreground and the background.
A.7 Time and Memory Complexity
Table 8: Runtime and memory comparison between SDXL, ToMe, and CO3 at different sampling budgets.
ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$ Time(s) $\downarrow$ Memory $\downarrow$
Method Animals Animals Objects Objects Animals Animals Objects Objects
SDXL (50 steps) 0.4906 1.1503 0.1131 0.6428 0.8378 0.4752 0.6183 7.20 8.5GB
ToMe (50 steps) 0.3894 1.5736 1.0117 0.6257 0.8807 0.6440 0.8542 16.58 18GB
CO3 (50 steps) 1.2341 1.6743 1.0158 0.7440 0.8877 0.5145 1.0120 19.90 9.5GB
CO3 (20 steps) 0.8532 1.6440 0.7481 0.6698 0.8611 0.4315 0.8679 10.65 9.5GB
We compare CO3 with its most competitive baseline ToMe taihang Hu et al. (2024) in terms of runtime peak memory and time requirement to generate one sample in Table 8. The experiments are conducted on a NVIDIA A100 gpu.
Although CO3 takes more time compared to ToMe with 50 time-steps, it requires significantly less memory. CO3 is lightweight, and its peak memory usage is only slightly above that of the base SDXL model. ToMe, being a gradient-based method, has substantial memory requirements.
We also emphasize that CO3 is model-agnostic and can be applied to any diffusion-based model. In contrast, ToMe is model-architecture dependent, as its attribute-binding mechanism operates by manipulating internal attention activations of the base model.
For a fair comparison, we additionally tested CO3 with 20 inference steps. As shown in the table below, the average performance of CO3 with 20 steps remains comparable to, and in many cases better than, ToMe with 50 steps. CO3 with $20$ steps takes $10.65$ seconds, which is faster than ToMe.
A.8 Ablations on Hyperparameters